A plane has an air speed of 240km/h.what should be the planes heading if it is to travel due north,relative to the earth in a wind blowing with a velocity of 100km/h in an easterly direction?
1
Expert's answer
2019-12-16T10:55:01-0500


Airplane to travel due north if

"\alpha=asin(\frac{v_w}{v_a})"

"\alpha=asin(\frac{100}{240})=24.6^o"

The aircraft should heading 24.6 degrees west of east